Hess’s law lab data and calculations
Data and Calculations NAME INSTRUCTOR mass of Mg volume HCI mass HCl(density of solution mass of solution + Mg initial temperature final temperature ΔΤ -1.05g/ml) oC oC oC ΔΗ, ( for 1 mole of Mg) kJ/mol Average ΔΗ3 kJ/mol o550 mass of Mgo volume HC mass HCl(density of solution- 1.05g/mL) 56 mass of solution +MgO initial temperature final temperature grams mL oC ec oC oC Ical ΔΙ Inn kJ/mol kJ/mol ΔΗ4 (for 1 mole of MgO) Average ΔΗ4 Using equations 2, 3, 4, and their AH values, caleulate kJ/mol AH Show calculations on the back of this page. kJ/mol Hess's Law
Co Ca G(3) CO2 SO2 NOs Total No. of Valence Electrons Lewis Structure(s) (Show all Resonance forms) Draw Structure of any Isomers the Geometrie Shape (State the geometry) Is the structure polar or nonpolar?
Answer #1

1. For the reaction of Mg and HCl:

Mg + 2 HCl === MgCl2 + H2

delta T = Final temperature - Initial Temperature

Delta T = 31.6 - 23 = 10.3 C

b) q calorimeter, the equation you need is:

q calorimeter = mass of calorimeter * heat capacity * delta T

according to the statement the mass of the solution is 52.586 grams

heat capacity will be 4.184 J / g C, same as water typically used in this problem

so

q calorimeter = 52.586*4.184*10.3 = 2266.204 Joules

c)

According to the energy balance

q rxn + q calorimeter = 0

q rxn = - qcalorimeter

q rxn = -2 266.204 Joules = - 2.266 KJ

d ) mass of mg = 0.086 grams

molar mass of Mg = 24.305

moles of Mg = 0.003538 moles

H *J / mole Mg = - 2266.204 / 0.003538 = - 640.532 KJ / moles Mg

For trial 2

a)delta T = 33.4 - 21.6 = 11.8 C

q cal = 52.6 grams * 4.184 J / g C * 11.8 C

q cal = 2596.92 Joules

q rxn = - 2.59 KJoules

moles of Mg = mass / molar mass

moles = 0.1 / 24.3 = 0.004115

H rxn = -2.59 / 0.004115 = -629.4 KJ / mole Mg

average H rxn =- 634.95 KJ / mole

the negative sign is because the system is releasing heat which is being absorbed by the calorimeter